# Less is More: Token-Efficient Video-QA via Adaptive Frame-Pruning and Semantic Graph Integration

### v5

@@ B.4. Representative Frame Selection Strategy
A crucial step in our AFP algorithm is selecting a single representative frame from each generated cluster. We investigated two primary strategies for this selection process. To ensure a robust conclusion, we conducted multiple runs for each strategy and report the averaged results. The experiments were performed on the LongVideoBench dataset with the Qwen2.5-VL-7-Instruct model, starting from Top-8 keyframes.
  • Strategy X (Highest Score First): This strategy selects the frame from each cluster that has the highest initial relevance score provided by the upstream VSLS selector. It leverages external, task-specific relevance information.
  • Strategy Y (Centroid Selection): This strategy selects the frame that is the most representative of the cluster internally. For each cluster, we identify the ”centroid frame”—the frame that has the minimum average feature distance to all other frames within that same cluster. This method is self-contained and relies solely on the visual features used for clustering.
Strategy Avg. Accuracy (%) Avg. Frames
Strategy X (Highest Score) 40.96 2.26
Strategy Y (Centroid) 41.17 2.14
Table 6: Comparison of representative frame selection strategies based on averaged results over multiple runs. The Centroid Selection strategy (Y) demonstrates superior overall performance and efficiency.
As shown in Table 6, the averaged results over multiple runs confirm that the Centroid Selection strategy (Strategy Y) holds a clear advantage. It not only achieves a higher weighted average accuracy across the dataset but also does so more efficiently, producing a more compact set of keyframes on average. While Strategy X shows strong performance on short and long videos, Strategy Y’s significant lead on medium-length videos contributes to its superior overall score. Both methods exhibit comparable stability across runs.
The superiority of the Centroid strategy is theoretically sound. By selecting the frame closest to the cluster’s feature-space center, it guarantees the most visually representative frame is chosen, minimizing information loss during the pruning process. This self-contained logic is arguably more robust and generalizable than relying on external scores, which might be noisy or biased towards the initial selection method. Based on these comprehensive findings, we adopted the Centroid Selection strategy for all main experiments reported in this paper.
@@
To validate the universal applicability of our AFP framework, we conducted a comprehensive set of experiments by applying it to the output of another state-of-the-art keyframe selector, T* (Ye et al. 2025). This section presents a full generalizability and ablation study across all three tested MLLMs (GPT-4o, Qwen2.5-VL, LLaVA-Video) on both the LongVideoBench and VideoMME datasets.
Unlike VSLS, T* does not explicitly extract semantic-logical relations. Therefore, for all experiments involving a semantic graph, we employed our low-cost, text-only LLM-based generation strategy (Case 2, described in the main paper’s Methodology) to create the graph.
@@
Notably, the performance gains are again most pronounced on the open-source models. Using LLaVA-Video-7B on VideoMME (from Top-32), our method elevates the accuracy on short videos from 37.7% (T* Baseline) to 55.0%—a staggering 17.3-point absolute improvement—while using only 4.1 frames instead of 32. This reinforces our main finding that open-source MLLMs benefit immensely from our pruning and compensation strategy. Furthermore, the success of the LLM-generated semantic graph in this setting validates its effectiveness as a low-cost, universally applicable alternative when pre-existing semantic information is unavailable. This confirms that the benefits of our framework are not contingent on a specific upstream selector like VSLS, but stem from the fundamental principles of intelligent redundancy reduction and semantic compensation.

### v6

@@ B.4. Representative Frame Selection Strategy
A crucial step in our AFP algorithm is selecting a single representative frame from each generated cluster. To determine the most effective and robust approach, we conducted a comprehensive ablation study comparing three distinct strategies. The experiments were performed on the LongVideoBench dataset with the GPT-4o model, starting from Top-8 keyframes provided by VSLS, and the results reported are averaged over multiple runs to ensure reliability. This analysis focuses on the ‘AFP only‘ setting (without the semantic graph) to purely evaluate the quality of the selected visual information.
  • Score-based (Highest VSLS Score): This strategy leverages external information by selecting the frame from each cluster that has the highest initial relevance score assigned by the upstream VSLS selector.
  • Centroid-based (Visual Centroid): This strategy is self-contained. It selects the ”centroid frame”—the frame with the minimum average feature distance to all other frames within the same cluster, making it the most visually representative.
  • Relevance-based (Highest Query Similarity): This strategy introduces task-specific guidance by selecting the frame with the highest CLIP similarity score to the ‘Question + Options‘ text prompt.
Experimental Settings Averaged Results
Dataset Model Frame Source Selection Strategy Avg. Acc (%) Avg. Frames
LongVideoBench GPT-4o VSLS (Top-8) Score-based 47.14 2.1
Centroid-based 48.53 2.1
Relevance-based 46.67 2.2
Table 6: Ablation study on representative frame selection strategies. The experiment was conducted in the ‘AFP only‘ setting to purely evaluate the quality of the selected visual information. Results are averaged over multiple runs. The Centroid-based strategy demonstrates the best performance.
As shown in Table 6, the averaged results over multiple runs confirm that the Centroid-based strategy holds a clear advantage. It achieves the highest overall accuracy (48.53%), outperforming both the Score-based and Relevance-based approaches.
The superiority of the Centroid strategy is theoretically sound, especially in the absence of a semantic graph. By selecting the frame closest to the cluster’s feature-space center, it guarantees the most visually representative frame is chosen, which is the most robust way to minimize information loss during the pruning process. This self-contained logic is arguably more generalizable than relying on external scores (from VSLS or CLIP), which might be noisy or biased. Based on these comprehensive findings, we adopted the Centroid-based strategy for all main experiments reported in this paper.
@@
To validate the universal applicability of our AFP framework as a post-processing module, we conducted a comprehensive set of experiments by applying it to the output of another state-of-the-art keyframe selector, T*. This section presents a full generalizability and ablation study across all three tested MLLMs (GPT-4o, Qwen2.5-VL-7B-Instruct, and LLaVA-Video-7B-Qwen2) on both the LongVideoBench and VideoMME datasets.
@@
A Note on Semantic Graph Generation for T*.
A key point in our experimental design is the source of the semantic graph for the AFP + Graph experiments in this section. While the T* method itself does not explicitly extract semantic-logical relations, our experiments for both T* and VSLS were conducted on the exact same set of Video-QA instances from the LongVideoBench and VideoMME benchmarks. This allows us to directly reuse the semantic graphs that were already generated for the VSLS experiments for each corresponding data instance. This approach is not only highly efficient, maintaining the zero-cost nature of our primary setup, but also ensures a perfectly fair comparison: the semantic graph, as a variable, is held constant, isolating the performance difference to the quality of the initial keyframe sets provided by T* versus VSLS.
@@
Notably, the performance gains are again most pronounced on the open-source models. Using LLaVA-Video-7B-Qwen2 on VideoMME (from Top-32), our method elevates the accuracy on short videos from 37.7% (T* Baseline) to 55.0%—a staggering 17.3-point absolute improvement—while using only 4.1 frames instead of 32. This reinforces our main finding that open-source MLLMs benefit immensely from our pruning and compensation strategy. Furthermore, the success of the LLM-generated semantic graph in this setting validates its effectiveness as a low-cost, universally applicable alternative when pre-existing semantic information is unavailable. This confirms that the benefits of our framework are not contingent on a specific upstream selector like VSLS, but stem from the fundamental principles of intelligent redundancy reduction and semantic compensation.
Figure 11: Generalizability of AFP + Graph on T* Keyframes. The plot shows the performance-efficiency trade-off on both LongVideoBench (left) and VideoMME (right) datasets across three MLLMs. Solid lines represent the high-cost T* baseline at 8 and 32 frames. Dashed lines represent our low-cost AFP + Graph method applied to the corresponding T* outputs. Our method consistently operates in the top-left (high-accuracy, low-frame-count) region, demonstrating its effectiveness as a universal refinement module.
The results robustly visualized in Figure 11, confirm that our AFP framework is a versatile plug-and-play refinement module. As shown in the plots, our AFP + Graph method (dashed lines) consistently operates in the highly desirable ”top-left” quadrant of the performance-efficiency space. It achieves an accuracy that is either competitive with or, in many cases, significantly superior to the high-cost T* baseline (solid lines), while using only a fraction of the frames. For instance, on LongVideoBench with the LLaVA-Video-7B model, our method applied to T* Top-8 outputs (2.2 frames) achieves a 51.15% accuracy, surpassing the T* Top-8 baseline (48.06%) by over 3 absolute points. This demonstrates that AFP + Graph is not overfitted to the properties of a single selector like VSLS and can effectively enhance other keyframe selection methods.